### 30 Jan 2024 v1

# Augmenting Replay in World Models for Continual Reinforcement Learning

Abstract
In continual RL, the environment of a reinforcement learning (RL) agent undergoes change. A successful system should appropriately balance the conflicting requirements of retaining agent performance on already learned tasks, stability, whilst learning new tasks, plasticity. The first-in-first-out buffer is commonly used to enhance learning in such settings but requires significant memory. We explore the application of an augmentation to this buffer which alleviates the memory constraints, and use it with a world model model-based reinforcement learning algorithm, to evaluate its effectiveness in facilitating continual learning. We evaluate the effectiveness of our method in Procgen and Atari RL benchmarks and show that the distribution matching augmentation to the replay-buffer used in the context of latent world models can successfully prevent catastrophic forgetting with significantly reduced computational overhead. Yet, we also find such a solution to not be entirely infallible, and other failure modes such as the opposite — lacking plasticity and being unable to learn a new task — to be a potential limitation in continual learning systems.
1 Introduction
Over a lifetime, a person or intelligent system that interacts with a complex environment can expect to acquire new experiences and learn new tasks, all while old experiences and tasks are retained. This requirement to satisfy two orthogonal goals — the acquisition of new knowledge and the retention of old knowledge is the challenge of continual learning (CL) Ring (1994), and the demand for both desiderata presents the stability-plasticity dilemma Carpenter and Grossberg (1987). A system capable of successfully undertaking CL should meet both of these goals. That is, it should be able to replicate its original performance on an earlier task when revisiting it after having learned other tasks, all while not interfering with its ability to gain performance on new tasks.
Neural networks have proved a powerful tool driving the state-of-the-art in games, control and interactive language models Vinyals et al. (2019); OpenAI et al. (2019); Beattie et al. (2016); Andrychowicz et al. (2020); Brown et al. (2020). However, neural networks have been shown to be highly susceptible to forgetting previously learnt knowledge after training on a new task French (1999). This undesirable property of neural networks has proven troublesome in their integration into CL systems, since catastrophe forgetting is the most reported failure mode McCloskey and Cohen (1989); Khetarpal et al. (2022). Catastrophic forgetting occurs when newly acquired skills compromise already learned ones. The opposite of catastrophic forgetting is forward and backward transfer, in which either earlier or later tasks benefit from learning on separate tasks.
Typically, the traditional RL settings focus on solving one individual task that does not change during learning Khetarpal et al. (2022). Where multiple different tasks are required to be learned, one common approach is to learn all simultaneously, rather than sequentially. This avoids catastrophic forgetting as the data distribution of the learning task never changes, but it is not always applicable, as the environments must generate too many samples or because some tasks are unavailable until others have completed. This is the case in continual RL, where new challenges are often experienced one after another. Therefore, a CL system must be able to learn multiple different tasks sequentially. Additionally, tasks may change subtly or silently over time, so a CL system should also be capable without explicit task identifiers Rolnick et al. (2019).
One method that existing CL systems employ in RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge from previous tasks to prevent catastrophic forgetting Lipton et al. (2018). This is done through a replay buffer, often implemented as a generic first-in-first-out (FIFO). Given that the buffer is sufficiently large and that the learning system is capable of using the knowledge stored in the buffer (i.e., the RL algorithm is not a purely on-policy method), such a solution may enable favourable properties of CL such as reduced forgetting and forward and backward transfer. However, the requirement to store all experiences from all tasks, demands very high storage capacity. This is an unfavourable property, as RL is often done at scale and the large volume of samples to store can present a significant engineering challenge, reducing the scalability and accessibility of the algorithm OpenAI et al. (2019).
The inspiration for many replay-based methods comes from Complementary Learning Systems (CLS) Hassabis et al. (2017); Khetarpal et al. (2022), which describes learning in mammalian brains. The hippocampus memorises recent observations and replays them to the neocortex, which is a slow statistical learner. Replay is interleaved with new experiences, thus mitigating catastrophic forgetting. Interestingly, the neocortex is understood to form a ‘world model’ whose purpose is to predict the consequences of our actions Mathis (2023).
The idea of world models has been exploited in RL algorithms (Model Based RL), where a world model is capable of predicting the effects of actions on the environment Ha and Schmidhuber (2018); Hafner et al. (2019, 2020, 2022, 2023). RL world model approaches are an intuitive choice for CL, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance. Although CLS describes replay to a ‘world model’, traditionally replay in RL is used in Model Free RL to improve the policy, rather than a world model directly.
In this paper, we take the most recent development of world model approaches in DreamerV3 Hafner et al. (2023) and explore its ability in CL when used with memory-efficient replay buffers Isele and Cosgun (2018). Hence, we present World Models with Augmented Replay (WMAR) as a promising solution to CL problems faced by world model approaches, while alleviating memory requirements. We evaluate WMAR on a diverse and challenging set of environments from OpenAI Procgen (procedurally generated games) and Atari. These suites pose distinct challenges as to how similar or dissimilar different tasks are. We used Procgen to create similar tasks with ‘shared structure’, whereas Atari games comprise dissimilar tasks ‘without shared structure’. Our analysis is not restricted to catastrophic forgetting but includes forgetting and backward/forward transfer of skills between tasks.
For a background in RL and CL, including desiderata in CL solutions, see Section S.LABEL:S-sec:background. For Related Work, see Section S.LABEL:S-sec:related_work; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al. (2017); Schwarz et al. (2018); Shin et al. (2017) and continual RL Isele and Cosgun (2018); Rolnick et al. (2019); Kessler et al. (2022).
2 World Models with Augmented Replay (WMAR)
WMAR uses world models and augmented replay buffers to learn tasks sequentially. The world model builds on developments in DreamerV3, and is state-of-the-art in single-GPU RL benchmarks in a number of tasks. The off-policy nature of the RL algorithm allows for the use of replay buffers as a mechanism to prevent forgetting. Furthermore, world models naturally support a memory-based continual learning approach as the strength of the final policy is largely contingent on the quality of the world model’s accuracy in simulating the environment. Hence, preventing forgetting within the world model through any means should retain agent performance in past environments. World models are also task-agnostic and labels specifying the task/environment are not required.
The system consists of the World Model for modelling the environment, an Actor-Critic Controller to act on the environment, and a Replay Buffer. The Replay Buffer is used to train the World Model, and the World Model is able to ‘dream’ experiences, which are used to train the Controller. Each is described below, and more details are found in Section S.LABEL:S-sec:world_models.
The source code is available at https://github.com/cerenaut/wmar.
2.1 World Model
The World Model, see Figure 1, learns a Recurrent State-Space Model (RSSM) Hafner et al. (2019) to predict environment dynamics. At timestep $t$, the RSSM contains a deterministic hidden state $h_{t}$ and stochastic state $z_{t}$, either predicted by output of a variational autoencoder $z_{t}\sim q_{\theta}(z_{t}\mid h_{t},x_{t})$ or by the dynamics predictor from $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid h_{t})$ in open-loop prediction. RSSM models the stochastic representation of the next state given current and previous observations $x_{1:t}$ and actions $a_{1:t}$. It achieves this through a recurrent layer in the GRU, predicting the deterministic state $h_{t+1}=\text{GRU}(h_{t},z_{t},a_{t})$ and the stochastic state $\hat{z}_{t+1}=f(h_{t+1})$.
Following DreamerV3, the stochastic state $z_{t}$ consists of 32 discrete stochastic units with 32 categorical classes each. Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the undisclosed environment state. During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist and the priors are taken for the posteriors instead (i.e. $z_{t}=\hat{z}_{t}$). We use the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
At timestep $t$, the model state of the trained world model is a Markovian representation of the environment and is the concatenation of the deterministic state $h_{t}$ and stochastic state $z_{t}$. The world model is trained by reconstructing input images, rewards and episode continuation flags ($c_{t}\in\{0,1\}$ depending on whether the episode terminates after the observed state) from the model state. The RSSM component is trained through KL balancing Hafner et al. (2022), which allows the dynamics predictor to model the transition between states.
(a) World Model
(b) Flow of data within the RSSM
Figure 1: World Model overview and RSSM details.
2.2 Actor Critic
The actor and critic (Figure S.LABEL:S-fig:actor_critic) are MLPs mapping the model state to actions and value estimates. They are trained entirely on trajectories generated stochastically by the world model, which is referred to as ‘dreaming’ or ‘imagining’. Starting from model states of inputs sampled from the replay buffer, the actor acts within states predicted by the world model, producing a sequence of states, actions, rewards, and continuation flags. We follow DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992). As such, inevitable changes in distribution within the model state space are not an issue since imagined trajectories are cheap to generate and do not require interaction with the actual environment, allowing this process to be run to convergence of the actor and critic.
2.3 Replay Buffer
A key objective of WMAR is to limit the size of the replay buffers. While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically choose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance. We also introduce spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes. The buffer types are shown in Figure 2 and described below.
Figure 2: The use of replay within the world model context. LTDM refers to the long-term distribution matching buffer.
2.3.1 Short-Term FIFO Buffer
We maintain a FIFO buffer with a capacity of $2^{18}$ samples containing the most recent rollout observations. The retention of a short-term FIFO buffer follows CLS theory, suggesting that multiple separate structures within a system are required for learning and CL. This buffer also allows the world model to train on all incoming experiences and biases the training samples to those more recently collected, thus improving convergence properties on the current task.
2.3.2 Long-Term Global Distribution Matching (LTDM)
LTDM Buffer
Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele and Cosgun (2018). We use a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contain a uniformly random subset of 512 spliced rollouts. This buffer uses the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving the experiences with the highest key values, and discarding the remainder.
Spliced Rollouts
Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially as episode lengths may be in the thousands. Containing too few unique episodes may cause the training data to become highly unrepresentative of the general environment states, thus reducing the world model’s accuracy in modelling the environment and likely causing a subsequent loss of performance from the actor. This is especially relevant for the long-term global distribution matching buffer, which may store relatively few samples from each environment.
Hence, while typical implementations store complete rollouts, we splice rollouts into smaller chunks of size 512 where necessary, which we find to not adversely affect performance. The implication is that the long-term global distribution matching buffer will operate over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
Remainders of rollouts with fewer than 512 states are concatenated prior to the next episode with an appropriate reset flag indicating the start of a new episode. To improve the efficiency of the rollout process, we truncate episodes after a fixed number of steps, so that the final number of environment steps is identical in every training iteration. We find spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
2.3.3 Augmented Replay Buffer
We combine the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both buffers in parallel. Data from both buffers are uniformly sampled for each training minibatch (Algorithm S.LABEL:S-alg:combined_buffers).
2.4 Task Agnostic Exploration
The exploration of new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task. To address this challenge, we use developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor critic training step. We find that these developments enable mitigating the exploration challenge with a fixed entropy regulariser and are able to produce strong results without the addition of an exploration-oriented learning system such as Plan2Explore (Section S.LABEL:S-sec:task_agnostic). Avoiding the need to train and run inference on an ensemble of predictors also provides benefits in wall time.
3 Experiments
We perform an empirical study to evaluate the efficacy of WMAR for CL in a set of challenging environments from OpenAI Procgen and Atari Cobbe et al. (2020); Bellemare et al. (2013), which we select for the unique challenges they present in a CL setting, where environments may be extremely similar (Procgen) or dissimilar (Atari). As measures of baseline performance, we compare the performance of WMAR with the most similar model-free and model-based RL algorithms on individual tasks. We choose proximal policy optimisation (PPO) Schulman et al. (2017); Raffin et al. (2021) for the model-free RL algorithm, and use the CNN-LSTM network architecture as it most similarly resembles the architecture of WMAR. We use DreamerV3 as the model-based benchmark.
To measure the efficacy of the long-term distribution matching buffer, we compare CL metrics for WMAR and WMAR with only the FIFO buffer and compare them both with DreamerV3 on Atari. Finally, we compare the performance difference of WMAR when applied to multiple sequential tasks versus when applied to only single tasks. All WMAR experiments are trained on one Nvidia A40 40GB GPU each with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
For more details on experimental parameters and execution time, see Section S.LABEL:S-sec:extime.
3.1 Evaluation
We evaluate the performance by normalising episodic reward to the performance provided by the world model agent’s baseline, and that of a random agent. This enables a normalised comparison in order to measure the relative performance during CL. These results also enable a natural evaluation of forgetting and forward transfer for all environments in each CL suite.
For each CL experiment, we define an ordered suite of tasks $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We define the performance of the agent on task $\tau\in\boldsymbol{\mathcal{T}}$ after $N$ environment steps as $p^{\boldsymbol{\mathcal{T}}}_{\tau}(N)$.
3.1.1 Baseline Normalisation
We perform 2 evaluations of each environment to be considered as ‘baseline’ performance. For each task in the experimental suite $\tau\in\boldsymbol{\mathcal{T}}$, we find the average episodic reward of a random policy $p^{(\tau)}_{\tau}(0)$ and the average episodic reward of the agent on only task $\tau$ after $N$ environment steps $p^{(\tau)}_{\tau}(N)$. The normalized score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
$q^{\boldsymbol{\mathcal{T}}}_{\tau}(N)=\frac{p^{\boldsymbol{\mathcal{T}}}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}{p^{(\tau)}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}$ (1)
Average random performance and average final performance of WMAR, when trained on a single task, are used for score normalisation. A score equal to the random performance would correspond to 0, and a score corresponding to the average final performance would correspond to 1.
3.1.2 Continual Learning Assessment
We follow Kessler et al. (2022) in evaluating the efficacy of a continual learning method in forgetting, measuring stability and backward transfer, and forward transfer, also acting as a measure for plasticity. These measurements are recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We train the agent on each task for $N$ environment steps. As such, the reference score for task $\tau_{i}$ would be given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
Forgetting
The average forgetting for each task is the difference between its performance at the end of training all tasks and its reference score.
$F=\frac{1}{T}\sum_{i=1}^{N}q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(T\times N)-q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$ (2)
A lower value for forgetting is indicative of improved stability and a superior continual learning method. A negative value for forgetting would imply that the agent has managed to gain performance on earlier tasks, thus exhibiting backward transfer.
Forward transfer
The forward transfer for each task is the expected ratio of average normalised scores between its performance in the continual learning and single task experiments.
$FT$ $=\frac{1}{T}\sum_{i=1}^{T}\frac{S^{\boldsymbol{\mathcal{T}}}_{i}-S^{(\tau_{i})}}{S^{(\tau_{i})}}$ (3)
where
$S^{\boldsymbol{\mathcal{T}}}_{i}$ $=\frac{1}{N}\sum_{j=1}^{N}q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}((i-1)\times N+j)$ (4)
$S^{(\tau_{i})}$ $=\frac{1}{N}\sum_{j=1}^{N}q_{\tau_{i}}^{(\tau_{i})}(j)$ (5)
The larger the forward transfer, the better the continual learning method. A positive value for forward transfer would imply that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result. When each task is not related to each other, no positive forward transfer may be expected. In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks by previous tasks.
3.2 Environments
We explore the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards. To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to distinguish between different environments. For evaluating the performance on tasks without shared structure, we use 4 different tasks from Atari.
4 Results
4.1 Tasks with Shared Structure
The CoinRun environment provides a number of simple visual perturbations. While the environment’s mechanics remain identical, the environments share both subtle and significant visual differences, which can pose significant risks to CL characteristics. However, we show that WMAR enables constructive interference and positive forward transfer of skills between these environments.
Baseline Performance
We measure the performance of WMAR and other baseline RL algorithms when applied to each task within the suite individually (see Section S.LABEL:S-sec:baseline for parameters and results). The performance of WMAR is approximately 15% better across tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences. The ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
Continual Learning
The normalised performance is shown in Figure 3 (tabular results in Section LABEL:S-sec:cl). The terminology that denotes the variant of the CoinRun environment is as follows. NB = no background, RT = restricted themes, and MA = Monochrome assets (see Section S.LABEL:S-sec:procgen).
Forgetting and Backward Transfer
Regardless of the choice of replay buffer, the agent is capable of exhibiting desirable forgetting and backward transfer properties. However, we observe a slight benefit to those desiderata when the long-term global distribution matching buffer is used in conjunction with the FIFO buffer, evidenced by the improved forgetting and backward transfer (Table 1). For the FIFO only implementation, no backward transfer on the ‘CoinRun’ environment is observed. This contrasts to the backward transfer observed for the same environment where the distribution matching buffer is present (Section S.LABEL:S-tab:raw_res_coinrun). Further, we observe that the method using the distribution matching buffer shows more consistent performance across tasks with fewer fluctuations in performance during training (Figure 3), which may be a desirable property if there exists a need to snapshot and use the agent at arbitrary points during training.
Forward Transfer
Our results show no evidence of inhibition to the strength of the agent when it is required to learn multiple tasks with shared structure. Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer, see Table 1. We find significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance. Importantly, the ordering of the tasks is such that levels in later tasks are either subsets of, or different to levels in earlier tasks. This is so that the agent is unable to completely generalise on earlier tasks just by learning later tasks. Crucially, our results show that strong forward transfer also holds for tasks that are not subsets of earlier tasks.
Figure 3: Continual learning performance of WMAR with (top) or without (bottom) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Table 1: Average forgetting and forward transfer on tasks with shared structure.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{-0.07}\pm 0.02$
$\mathbf{0.34}\pm 0.06$
WMAR (FIFO)
$-0.04\pm 0.03$
$0.32\pm 0.06$
4.2 Tasks without Shared Structure
We use a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and follow Machado et al. (2018) in using sticky actions in Atari environments. We run the CL experiment on an arbitrary ordering of the environments as testing each permutation would be prohibitively expensive. We show that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL. Moreover, lacking a task oracle, we find predetermined reward scales (Table S.LABEL:S-tab:atari_rew_scales) for each individual environment necessary to enable continual learning.
Baseline Performance
WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Section S.LABEL:S-sec:baseline for parameters and results). Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
Continual Learning
The normalised performance is plotted in Figure 4 (tabular results in Section LABEL:S-sec:cl). It is clear that DreamerV3 is not capable of CL; it only performs well when it is actively training on a given task.
Forgetting
WMAR shows significantly improved stability in learning at all points after the task is learned. With only the FIFO buffer, and in all tasks except the last, the agent loses virtually all performance before the following task has finished training. This is in direct contrast to WMAR with the addition of the long-term global distribution matching buffer where instead, the agent is able to consistently maintain much of its performance on tasks learned earlier on (Table 2). This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model. As training progresses, samples from the first task have a decreasing probability of being drawn. This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn. Despite this, our results show that WMAR is able to hold much of its performance on the earliest of tasks (Figure 4). Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
Figure 4: Continual learning performance of WMAR with (top) or without (middle) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. DreamerV3 benchmark results are also given (bottom). Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Table 2: Average forgetting and forward transfer on tasks without shared structure.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{0.10}\pm 0.04$
$-0.35\pm 0.10$
WMAR (FIFO)
$0.68\pm 0.10$
$\mathbf{-0.11}\pm 0.14$
Forward Transfer and Plasticity
We find significant differences in the properties of continual learning on tasks with and without shared structure which are reflected in the results. Contrary to findings in tasks with shared structure, the continual learning setting offers no benefit to forward transfer. Benefits to continual learning in greatly reduced forgetting come at the cost of also reduced forward transfer — effectively trading some plasticity for a significant amount of stability, as the agent is unable to learn later tasks as effectively. We observe a decreasing trend in the agent’s normalized score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’ (Table S.LABEL:S-tab:raw_res_atari), which is already well below 1.0. This evidences that it performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the agent. This limitation thus results in plasticity being gradually lost as training progresses. Moreover, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learned a previous task as compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in CL with world models.
5 Discussion
WMAR has the ability to perform CL with reduced memory requirements. On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits reduced forgetting. However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure), remains a challenge. Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learned less effectively if they are learned later.
Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer. This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers. Further, being able to learn tasks quickly benefits scalability of the algorithm.
Our results reiterate findings of world model approaches as strong baselines for CL problems, which we attribute to their natural ability to support off-policy experience replay. Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output. Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
5.1 Generalisation of Tasks
While we find tasks with shared structure to enable positive forward and backward transfer, we note that this is not the case in tasks without shared structure. We posit this is due to the reliance on the convolutional feature extractor within the WMAR architecture, enabling it to only take advantage of commonality in the visual domain. Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks. For example, a change to an environment involving inverting the colours and permuting the controls is likely to induce negative forward transfer. A key distinction between CL with world models and model-free methods is the requirement to reconstruct the high-dimensional observations in the form of images. We posit this induces significant strain on the capacity of the model, possibly in the learning of the latent space, or in the model parameters for encoding and reconstructing the observation, especially if it varies greatly between tasks.
5.2 Reward Scaling without Task Oracle
We also find that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales. While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we find this property to not hold when reward scales differ significantly (e.g. by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt. Barring the possibility of access to a task oracle, we believe that this limitation affects all continual learning applications. This limitation is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes. Experiments involving approaches with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
5.3 Limitations of Reservoir Sampling
The WMAR approach brings substantial improvement to memory requirements and CL characteristics, all while significantly reducing catastrophic forgetting. However, the scalability of this approach to a larger number of tasks remains unproven. A key limitation inherited from the use of reservoir sampling is the finite capacity of the long-term global distribution matching buffer. As more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting. Indeed, we observe some variance in the amount of forgetting to which each task is susceptible. Results from the experiments on tasks without shared structure (Figure 4) suggest latter tasks which may have been learned poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
A plausible approach to improving the performance of WMAR on a larger number of tasks is to increase the size of the long-term global distribution matching buffer. However, this approach naturally inherits limitations from world model approaches and memory-based CL approaches, as doing so incurs significant memory requirements. While techniques such as image compression may alleviate this constraint, RL is frequently done at scale, where the need to send and store a significant volume of image data may pose a significant engineering challenge, reducing the applicability and accessibility of this algorithm.
6 Conclusion and Future Work
We develop a powerful approach for CL within the context of RL approaches using world models. We expect future developments to the concepts shown here, or for existing techniques like behaviour cloning in CLEAR to be combined with our method for further performance gains. Such an adaptation could be countering shifts in the latent distribution as the world model trains and where the actor is frozen. Further, we posit the possibility of tuning the hyperparameters and configuration of WMAR to increase sample efficiency and model capacity to further improve CL. We also propose a difference in learning tasks with and without shared structure, where later tasks may be dissimilar in input, control, and reward scale, and hope for future research to explore the effect of such difference on CL within supervised and non-supervised tasks.

### 05 Mar 2024 v2

@@
 Abstract
-In continual RL, the environment of a reinforcement learning (RL) agent undergoes change.
-A successful system should appropriately balance the conflicting requirements of retaining agent performance on already learned tasks, stability, whilst learning new tasks, plasticity.
-The first-in-first-out buffer is commonly used to enhance learning in such settings but requires significant memory.
-We explore the application of an augmentation to this buffer which alleviates the memory constraints, and use it with a world model model-based reinforcement learning algorithm, to evaluate its effectiveness in facilitating continual learning.
-We evaluate the effectiveness of our method in Procgen and Atari RL benchmarks and show that the distribution matching augmentation to the replay-buffer used in the context of latent world models can successfully prevent catastrophic forgetting with significantly reduced computational overhead.
-Yet, we also find such a solution to not be entirely infallible, and other failure modes such as the opposite — lacking plasticity and being unable to learn a new task — to be a potential limitation in continual learning systems.
+Continual RL is a challenging problem where the agent is exposed to a sequence of tasks; it should learn new tasks without forgetting old ones, and learning the new task should improve performance on previous and future tasks.
+The most common approaches use model-free RL algorithms as a base, and replay buffers have been used to overcome catastrophic forgetting.
+However, the buffers are often very large making scalability difficult.
+Also, the concept of replay comes from biological inspiration, where evidence suggests that replay is applied to a world model, which implies model-based RL – and model-based RL should have benefits for continual RL, where it is possible to exploit knowledge independent of the policy.
+We present WMAR, World Models with Augmented Replay, a model-based RL algorithm with a world model and memory efficient distribution matching replay buffer.
+It is based on the well-known DreamerV3 algorithm, which has a simple FIFO buffer and was not tested in a continual RL setting.
+We evaluated WMAR vs WMAR (FIFO only) on tasks with and without shared structure from OpenAI ProcGen and Atari respectively, and without a task oracle.
+We found that WMAR has favourable properties on continual RL with significantly reduced computational overhead compared to WMAR (FIFO only).
+WMAR had small benefits over DreamerV3 on tasks with shared structure and substantially better forgetting characteristics on tasks without shared structure; but at the cost of lower plasticity seen in a lower maximum on new tasks.
+The results suggest that model-based RL using a world model with a memory efficient replay buffer can be an effective and practical approach to continual RL, justifying future work.
 1 Introduction
-Over a lifetime, a person or intelligent system that interacts with a complex environment can expect to acquire new experiences and learn new tasks, all while old experiences and tasks are retained.
-This requirement to satisfy two orthogonal goals — the acquisition of new knowledge and the retention of old knowledge is the challenge of continual learning (CL) Ring (1994), and the demand for both desiderata presents the stability-plasticity dilemma Carpenter and Grossberg (1987).
-A system capable of successfully undertaking CL should meet both of these goals.
-That is, it should be able to replicate its original performance on an earlier task when revisiting it after having learned other tasks, all while not interfering with its ability to gain performance on new tasks.
-Neural networks have proved a powerful tool driving the state-of-the-art in games, control and interactive language models Vinyals et al.
-(2019); OpenAI et al.
-(2019); Beattie et al.
-(2016); Andrychowicz et al.
-(2020); Brown et al.
+Typically, RL settings focus on one individual task that does not change during learning.
+In continual RL, new tasks are experienced one after another, which is more challenging due to the well known issue of catastrophic forgetting McCloskey & Cohen (1989); Khetarpal et al.
+(2022) but represents many real-world scenarios Khetarpal et al.
 (2022).
-However, neural networks have been shown to be highly susceptible to forgetting previously learnt knowledge after training on a new task French (1999).
-This undesirable property of neural networks has proven troublesome in their integration into CL systems, since catastrophe forgetting is the most reported failure mode McCloskey and Cohen (1989); Khetarpal et al.
+A simple approach to learning multiple tasks is to learn them simultaneously, but in continual RL, there are often too many environments for that to be practical and tasks are unavailable until others have completed.
+Additionally, tasks may change subtly or silently over time, so the system should not require explicit task identifiers Rolnick et al.
 (2019).
-Catastrophic forgetting occurs when newly acquired skills compromise already learned ones.
-The opposite of catastrophic forgetting is forward and backward transfer, in which either earlier or later tasks benefit from learning on separate tasks.
-Typically, the traditional RL settings focus on solving one individual task that does not change during learning Khetarpal et al.
+One approach to continual RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge to prevent catastrophic forgetting Lipton et al.
 (2018).
-Where multiple different tasks are required to be learned, one common approach is to learn all simultaneously, rather than sequentially.
-This avoids catastrophic forgetting as the data distribution of the learning task never changes, but it is not always applicable, as the environments must generate too many samples or because some tasks are unavailable until others have completed.
-This is the case in continual RL, where new challenges are often experienced one after another.
-Therefore, a CL system must be able to learn multiple different tasks sequentially.
-Additionally, tasks may change subtly or silently over time, so a CL system should also be capable without explicit task identifiers Rolnick et al.
-(2019).
-One method that existing CL systems employ in RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge from previous tasks to prevent catastrophic forgetting Lipton et al.
-(2018).
-This is done through a replay buffer, often implemented as a generic first-in-first-out (FIFO).
-Given that the buffer is sufficiently large and that the learning system is capable of using the knowledge stored in the buffer (i.e., the RL algorithm is not a purely on-policy method), such a solution may enable favourable properties of CL such as reduced forgetting and forward and backward transfer.
-However, the requirement to store all experiences from all tasks, demands very high storage capacity.
-This is an unfavourable property, as RL is often done at scale and the large volume of samples to store can present a significant engineering challenge, reducing the scalability and accessibility of the algorithm OpenAI et al.
+This is often done through a replay buffer, implemented as a generic first-in-first-out (FIFO).
+However, the requirement to store all experiences from all tasks, demands very high storage capacity, reducing scalability and accessibility of the algorithm OpenAI et al.
 (2019).
@@
 Interestingly, the neocortex is understood to form a ‘world model’ whose purpose is to predict the consequences of our actions Mathis (2023).
-The idea of world models has been exploited in RL algorithms (Model Based RL), where a world model is capable of predicting the effects of actions on the environment Ha and Schmidhuber (2018); Hafner et al.
-(2019, 2020, 2022, 2023).
-RL world model approaches are an intuitive choice for CL, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance.
+The idea of world models has been exploited in model-based RL algorithms, where a world model is capable of predicting the effects of actions on the environment Ha & Schmidhuber (2018); Hafner et al.
+(2019) and the more recent state-of-the-art DreamerV1 to V3 Hafner et al.
+(2020; 2022; 2023).
+To our knowledge, world model approaches have not been applied to continual RL.
+However, they are an intuitive choice, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance.
 Although CLS describes replay to a ‘world model’, traditionally replay in RL is used in Model Free RL to improve the policy, rather than a world model directly.
-In this paper, we take the most recent development of world model approaches in DreamerV3 Hafner et al.
-(2023) and explore its ability in CL when used with memory-efficient replay buffers Isele and Cosgun (2018).
-Hence, we present World Models with Augmented Replay (WMAR) as a promising solution to CL problems faced by world model approaches, while alleviating memory requirements.
-We evaluate WMAR on a diverse and challenging set of environments from OpenAI Procgen (procedurally generated games) and Atari.
-These suites pose distinct challenges as to how similar or dissimilar different tasks are.
-We used Procgen to create similar tasks with ‘shared structure’, whereas Atari games comprise dissimilar tasks ‘without shared structure’.
+In this paper, we take DreamerV3 Hafner et al.
+(2023), add a memory-efficient replay buffer Isele & Cosgun (2018), and apply it to continual RL.
+Hence, we present World Models with Augmented Replay (WMAR).
+We applied WMAR to two settings.
+First, where each task has a distinct environment and reward function.
+This is the most comment setting used in continual RL, and Atari games are often used.
+In the second setting, there is commonality between tasks, and learned knowledge can be leveraged to perform subsequent tasks.
+This is referred to as tasks with ‘shared structure’ Khetarpal et al.
+(2022).
+Often a video game is used, but the conditions such as movement dynamics or spacing of features in the environment change over time e.g.
+Riemer et al.
+(2019).
+Many potential real-world applications exist within the ‘shared structure’ setting.
+For example a robot assistant that should be able to acquire new and related tasks within a home like cleaning with a broom and then a mop.
+We used Procgen (procedurally generated games) from OpenAI to create similar tasks with ‘shared structure’, and Atari games for dissimilar tasks ‘without shared structure’.
 Our analysis is not restricted to catastrophic forgetting but includes forgetting and backward/forward transfer of skills between tasks.
-For a background in RL and CL, including desiderata in CL solutions, see Section S.LABEL:S-sec:background.
-For Related Work, see Section S.LABEL:S-sec:related_work; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al.
+For background and related work, see Appendix A; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al.
 (2017); Schwarz et al.
 (2018); Shin et al.
-(2017) and continual RL Isele and Cosgun (2018); Rolnick et al.
+(2017) and continual RL Isele & Cosgun (2018); Rolnick et al.
 (2019); Kessler et al.
 (2022).
+1.1 Desiderata
+Using a neural network, policy updates arising from the need to learn the current task often prove detrimental to preserving knowledge and performance on previous tasks.
+On the contrary, humans are able to accumulate knowledge to benefit performance on multiple related tasks simultaneously.
+Hence, a successful CL agent should be able to exhibit the following qualities Khetarpal et al.
+(2022); Chen & Liu (2018):
+  • Stability The agent avoids forgetting and losing performance on previously learnt tasks when learning new tasks.
+  • Backward Transfer The agent can increase its performance on previously learned tasks after training on new, similar tasks.
+  • Plasticity The agent, when encountering a new task unlike any of the previous, does not learn it any slower compared to learning it independently.
+  • Forward Transfer When encountering a new task that is similar to previously learnt tasks, the CL agent can learn it faster compared to learning it independently.
+Another challenge of continual RL is the lack of explicit task boundaries.
+Realistically, tasks may change unexpectedly, or evolve slowly over time.
+In these cases, a task oracle is either not available or not applicable.
+Further, practical considerations regarding CL such as its scalability and the wall time of experiments are a relevant point of interest.
+Therefore, the CL algorithm should also possess the following properties:
+  • Simple Integration of multiple systems supporting CL is a possibility in future research.
+WMAR proposes a simple and direct method of using replay within the context of RL with learned world models which may be combined easily with other approaches.
+Furthermore, WMAR is simple to implement and imposes minimal modification to the underlying RL algorithm, DreamerV3 Hafner et al.
+(2023).
+  • Scalable The solution should be able to scale to support CL on a large number of tasks.
+In order to do so, memory and computational constraints are necessary to ensure the wider accessibility and practicality of the solution.
+  • Does not require a task oracle A task oracle provides exact knowledge of which task is currently being trained, which may not be possible in practice.
+The requirement of a task oracle decreases the applicability of the solution.
+1.2 Contributions
+The primary contributions include: a) applying ‘world model’ DreamerV3 based architectures to the problem of continual RL for tasks with and without shared structure and b) enhancing the replay buffer of DreamerV3.
 2 World Models with Augmented Replay (WMAR)
 WMAR uses world models and augmented replay buffers to learn tasks sequentially.
-The world model builds on developments in DreamerV3, and is state-of-the-art in single-GPU RL benchmarks in a number of tasks.
-The off-policy nature of the RL algorithm allows for the use of replay buffers as a mechanism to prevent forgetting.
+The world model builds on developments in DreamerV3 which is state-of-the-art in single-GPU RL benchmarks in a number of tasks.
+The off-policy nature of the algorithm allows a replay buffer to prevent forgetting.
 Furthermore, world models naturally support a memory-based continual learning approach as the strength of the final policy is largely contingent on the quality of the world model’s accuracy in simulating the environment.
@@
 The Replay Buffer is used to train the World Model, and the World Model is able to ‘dream’ experiences, which are used to train the Controller.
-Each is described below, and more details are found in Section S.LABEL:S-sec:world_models.
+Each is described below, and more details are found in Appendix B.
 The source code is available at https://github.com/cerenaut/wmar.
@@
 Following DreamerV3, the stochastic state $z_{t}$ consists of 32 discrete stochastic units with 32 categorical classes each.
-Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the undisclosed environment state.
-During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist and the priors are taken for the posteriors instead (i.e.
+Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the hidden environment state.
+During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist, and the priors are taken for the posteriors instead (i.e.
 $z_{t}=\hat{z}_{t}$).
-We use the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
+We used the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
 At timestep $t$, the model state of the trained world model is a Markovian representation of the environment and is the concatenation of the deterministic state $h_{t}$ and stochastic state $z_{t}$.
@@
 2.2 Actor Critic
-The actor and critic (Figure S.LABEL:S-fig:actor_critic) are MLPs mapping the model state to actions and value estimates.
+The actor and critic (Figure 5, Appendix B.2) are MLPs that map the state of the model to actions and value estimates.
 They are trained entirely on trajectories generated stochastically by the world model, which is referred to as ‘dreaming’ or ‘imagining’.
 Starting from model states of inputs sampled from the replay buffer, the actor acts within states predicted by the world model, producing a sequence of states, actions, rewards, and continuation flags.
-We follow DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992).
+We followed DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992).
 As such, inevitable changes in distribution within the model state space are not an issue since imagined trajectories are cheap to generate and do not require interaction with the actual environment, allowing this process to be run to convergence of the actor and critic.
@@
 A key objective of WMAR is to limit the size of the replay buffers.
-While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically choose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance.
-We also introduce spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes.
+While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically chose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance.
+We also introduced spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes.
 The buffer types are shown in Figure 2 and described below.
@@
 2.3.1 Short-Term FIFO Buffer
-We maintain a FIFO buffer with a capacity of $2^{18}$ samples containing the most recent rollout observations.
+A FIFO buffer with a capacity of $2^{18}$ samples contained the most recent rollout observations.
 The retention of a short-term FIFO buffer follows CLS theory, suggesting that multiple separate structures within a system are required for learning and CL.
-This buffer also allows the world model to train on all incoming experiences and biases the training samples to those more recently collected, thus improving convergence properties on the current task.
+This buffer also allowed the world model to train on all incoming experiences and biased the training samples to those more recently collected, thus improving convergence properties on the current task.
 2.3.2 Long-Term Global Distribution Matching (LTDM)
 LTDM Buffer
-Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele and Cosgun (2018).
-We use a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contain a uniformly random subset of 512 spliced rollouts.
-This buffer uses the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving the experiences with the highest key values, and discarding the remainder.
+Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele & Cosgun (2018).
+We used a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contained a uniformly random subset of 512 spliced rollouts.
+This buffer used the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving experiences with the highest key values, and discarding the remainder.
 Spliced Rollouts
-Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially as episode lengths may be in the thousands.
+Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially since the duration of the episodes may be in the thousands.
 Containing too few unique episodes may cause the training data to become highly unrepresentative of the general environment states, thus reducing the world model’s accuracy in modelling the environment and likely causing a subsequent loss of performance from the actor.
 This is especially relevant for the long-term global distribution matching buffer, which may store relatively few samples from each environment.
-Hence, while typical implementations store complete rollouts, we splice rollouts into smaller chunks of size 512 where necessary, which we find to not adversely affect performance.
-The implication is that the long-term global distribution matching buffer will operate over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
+Hence, while typical implementations store complete rollouts, we spliced rollouts into smaller chunks of size 512 where necessary, which did not adversely affect performance.
+The implication is that the long-term global distribution matching buffer operated over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
 Remainders of rollouts with fewer than 512 states are concatenated prior to the next episode with an appropriate reset flag indicating the start of a new episode.
-To improve the efficiency of the rollout process, we truncate episodes after a fixed number of steps, so that the final number of environment steps is identical in every training iteration.
-We find spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
+To improve the efficiency of the rollout process, we truncated episodes after a fixed number of steps, so that the final number of environment steps was identical in every training iteration.
+We found spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
 2.3.3 Augmented Replay Buffer
-We combine the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both buffers in parallel.
-Data from both buffers are uniformly sampled for each training minibatch (Algorithm S.LABEL:S-alg:combined_buffers).
+We combined the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both in parallel.
+Data from both buffers were uniformly sampled for each training minibatch (Algorithm 2).
 2.4 Task Agnostic Exploration
-The exploration of new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task.
-To address this challenge, we use developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor critic training step.
-We find that these developments enable mitigating the exploration challenge with a fixed entropy regulariser and are able to produce strong results without the addition of an exploration-oriented learning system such as Plan2Explore (Section S.LABEL:S-sec:task_agnostic).
-Avoiding the need to train and run inference on an ensemble of predictors also provides benefits in wall time.
+Exploring new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task.
+To address this challenge, we used developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor-critic training step.
+These developments enabled mitigation of the exploration challenge with a fixed entropy regulariser and produced strong results without the addition of an exploration-orientated learning system such as Plan2Explore (Appendix B.4).
+Avoiding the need to train and run inference on an ensemble of predictors also benefits wall time.
 3 Experiments
-We perform an empirical study to evaluate the efficacy of WMAR for CL in a set of challenging environments from OpenAI Procgen and Atari Cobbe et al.
+We conducted an empirical study to evaluate the efficacy of WMAR for continual RL in a set of challenging OpenAI Procgen and Atari environments Cobbe et al.
 (2020); Bellemare et al.
-(2013), which we select for the unique challenges they present in a CL setting, where environments may be extremely similar (Procgen) or dissimilar (Atari).
-As measures of baseline performance, we compare the performance of WMAR with the most similar model-free and model-based RL algorithms on individual tasks.
-We choose proximal policy optimisation (PPO) Schulman et al.
+(2013).
+We selected them because they are commonly used as RL benchmarks, can be extremely similar for the case of shared structure (Procgen) or dissimilar in the case of without shared structure (Atari), and they were computationally feasible on the available hardware.
+We compared WMAR (with augmented replay buffer) with a baseline model equivalent to DreamerV3, referred to as ‘WMAR (FIFO Only)’.
+The comparison measures the efficacy of the long-term distribution matching buffer.
+To validate our implementation, we compared also to an open source implementation of DreamerV3 by the author https://github.com/danijar/dreamerv3, see Appendix D.1.
+To measure baseline performance, used for normalisation of other results, we also ran experiments on individual tasks, and compared WMAR with the most similar model-free and model-based RL algorithms.
+We chose proximal policy optimisation (PPO) Schulman et al.
 (2017); Raffin et al.
-(2021) for the model-free RL algorithm, and use the CNN-LSTM network architecture as it most similarly resembles the architecture of WMAR.
-We use DreamerV3 as the model-based benchmark.
-To measure the efficacy of the long-term distribution matching buffer, we compare CL metrics for WMAR and WMAR with only the FIFO buffer and compare them both with DreamerV3 on Atari.
-Finally, we compare the performance difference of WMAR when applied to multiple sequential tasks versus when applied to only single tasks.
-All WMAR experiments are trained on one Nvidia A40 40GB GPU each with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
-For more details on experimental parameters and execution time, see Section S.LABEL:S-sec:extime.
+(2021) for the model-free RL algorithm and used the CNN-LSTM network architecture, as it most closely resembles the WMAR architecture.
+Experimental setup
+All WMAR experiments are trained on one Nvidia A40 40GB GPU with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
+For more details on experimental parameters and execution time, see Appendix C.2.
 3.1 Evaluation
-We evaluate the performance by normalising episodic reward to the performance provided by the world model agent’s baseline, and that of a random agent.
+We evaluated the performance by normalising episodic reward to the performance provided by the world model agent’s baseline and that of a random agent.
 This enables a normalised comparison in order to measure the relative performance during CL.
 These results also enable a natural evaluation of forgetting and forward transfer for all environments in each CL suite.
+More details are given in the following subsections.
 For each CL experiment, we define an ordered suite of tasks $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$.
@@
 3.1.1 Baseline Normalisation
-We perform 2 evaluations of each environment to be considered as ‘baseline’ performance.
+We performed 2 evaluations of each environment to be considered as ‘baseline’ performance.
 For each task in the experimental suite $\tau\in\boldsymbol{\mathcal{T}}$, we find the average episodic reward of a random policy $p^{(\tau)}_{\tau}(0)$ and the average episodic reward of the agent on only task $\tau$ after $N$ environment steps $p^{(\tau)}_{\tau}(N)$.
-The normalized score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
+The normalised score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
 $q^{\boldsymbol{\mathcal{T}}}_{\tau}(N)=\frac{p^{\boldsymbol{\mathcal{T}}}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}{p^{(\tau)}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}$ (1)
-Average random performance and average final performance of WMAR, when trained on a single task, are used for score normalisation.
-A score equal to the random performance would correspond to 0, and a score corresponding to the average final performance would correspond to 1.
+Average random performance and average final performance of WMAR, when trained on a single task, were used for score normalisation.
+A score equal to the random performance corresponds to 0, and a score corresponding to the average final performance corresponds to 1.
 3.1.2 Continual Learning Assessment
-We follow Kessler et al.
+We followed Kessler et al.
 (2022) in evaluating the efficacy of a continual learning method in forgetting, measuring stability and backward transfer, and forward transfer, also acting as a measure for plasticity.
-These measurements are recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$.
-We train the agent on each task for $N$ environment steps.
-As such, the reference score for task $\tau_{i}$ would be given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
-Forgetting
+These measurements were recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$.
+We trained the agent on each task for $N$ environment steps.
+As such, the reference score for task $\tau_{i}$ is given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
+Forgetting and Backward Transfer
 The average forgetting for each task is the difference between its performance at the end of training all tasks and its reference score.
 $F=\frac{1}{T}\sum_{i=1}^{N}q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(T\times N)-q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$ (2)
-A lower value for forgetting is indicative of improved stability and a superior continual learning method.
+A lower value for forgetting is indicative of improved stability and a better continual learning method.
 A negative value for forgetting would imply that the agent has managed to gain performance on earlier tasks, thus exhibiting backward transfer.
@@
 The larger the forward transfer, the better the continual learning method.
-A positive value for forward transfer would imply that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result.
-When each task is not related to each other, no positive forward transfer may be expected.
-In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks by previous tasks.
+A positive value for forward transfer implies that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result.
+When each task is not related to the others, no positive forward transfer may be expected.
+In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks from previous tasks.
 3.2 Environments
-We explore the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards.
-To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to distinguish between different environments.
-For evaluating the performance on tasks without shared structure, we use 4 different tasks from Atari.
+We explored the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards.
+To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to create differences between tasks.
+For evaluating the performance on tasks without shared structure, we selected 4 Atari tasks.
 4 Results
+4.1 Tasks without Shared Structure
+We chose a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and followed Machado et al.
+(2018) in using sticky actions.
+We ran the CL experiment with arbitrary ordering of the environments as testing each permutation would be prohibitively expensive.
+We found that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL.
+Furthermore, without a task oracle, we found predetermined reward scales (Table 8) for each individual environment necessary to allow continual learning.
+Baseline Performance (Single Tasks)
+WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Appendix D.2 for parameters and results).
+The ‘Random’ and ‘WMAR’ scores are used as normalisation baselines for each task.
+Continual Learning
+The normalised performance is plotted in Figure 3 (tabular results in Section D.3, Table 8).
+WMAR is capable of continual RL, whereas it is clear that WMAR (FIFO only) the DreamerV3 equivalent, is not; it only performs well when it is actively training on a given task.
+Forgetting and Backward Transfer
+In all tasks except the last, WMAR (FIFO only) loses virtually all performance before the following task has finished training.
+This is in direct contrast to WMAR, where instead the agent can maintain much of its performance on previously learnt tasks, see Table 1.
+In summary, WMAR shows significantly improved stability.
+This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model.
+As training progresses, samples from the first task have a decreasing probability of being drawn.
+This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn.
+Despite this, our results show that WMAR is capable of holding much of its performance on the earliest of tasks.
+Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
+Forward Transfer
+Reduced forgetting comes at the cost of also reduced forward transfer – effectively trading some plasticity for a significant amount of stability – as the agent is unable to learn later tasks as effectively.
+We observe a decreasing trend in the agent’s normalised score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’, which is already well below 1.0.
+The system performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the policy.
+This limitation thus results in plasticity being gradually lost as training progresses.
+Furthermore, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learnt a previous task compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in continual RL with world models.
+Figure 3: Continual learning performance on tasks without shared structure.
+Bold line segments denote periods in which certain tasks are trained.
+Scores are normalised by Equation 1.
+The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
+Table 1: Average forgetting and forward transfer on tasks without shared structure.
+Model Avg. forgetting Avg. fwd. transfer
+(Lower is better)
+(Higher is better)
+WMAR
+$\mathbf{0.10}\pm 0.04$
+$-0.35\pm 0.10$
+WMAR (FIFO)
+$0.68\pm 0.10$
+$\mathbf{-0.11}\pm 0.14$
 4.2 Tasks with Shared Structure
 The CoinRun environment provides a number of simple visual perturbations.
-While the environment’s mechanics remain identical, the environments share both subtle and significant visual differences, which can pose significant risks to CL characteristics.
-However, we show that WMAR enables constructive interference and positive forward transfer of skills between these environments.
-Baseline Performance
-We measure the performance of WMAR and other baseline RL algorithms when applied to each task within the suite individually (see Section S.LABEL:S-sec:baseline for parameters and results).
-The performance of WMAR is approximately 15% better across tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences.
-The ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
+While the environment’s mechanics remain identical, the environments share both subtle and substantial visual differences, which poses significant challenges.
+However, we found that WMAR enables constructive interference and positive forward transfer of skills between these environments.
+Baseline Performance – Single Tasks
+The performance of WMAR is approximately 15% better in all tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences.
+Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
+See Appendix D.2 for parameters and detailed results.
 Continual Learning
-The normalised performance is shown in Figure 3 (tabular results in Section LABEL:S-sec:cl).
+The normalised performance is shown in Figure 4 (tabular results in Appendix D.3, Table 10).
 The terminology that denotes the variant of the CoinRun environment is as follows.
-NB = no background, RT = restricted themes, and MA = Monochrome assets (see Section S.LABEL:S-sec:procgen).
+NB = no background, RT = restricted themes, and MA = monochrome assets (see Appendix C.1).
+Both models can learn continuously and improve in performance and have approximately equal peak performance on each task.
+However, WMAR is noticeably more consistent than WMAR (FIFO Only).
 Forgetting and Backward Transfer
-Regardless of the choice of replay buffer, the agent is capable of exhibiting desirable forgetting and backward transfer properties.
-However, we observe a slight benefit to those desiderata when the long-term global distribution matching buffer is used in conjunction with the FIFO buffer, evidenced by the improved forgetting and backward transfer (Table 1).
-For the FIFO only implementation, no backward transfer on the ‘CoinRun’ environment is observed.
-This contrasts to the backward transfer observed for the same environment where the distribution matching buffer is present (Section S.LABEL:S-tab:raw_res_coinrun).
-Further, we observe that the method using the distribution matching buffer shows more consistent performance across tasks with fewer fluctuations in performance during training (Figure 3), which may be a desirable property if there exists a need to snapshot and use the agent at arbitrary points during training.
+Regardless of the choice of replay buffer, the agent exhibits desirable forgetting and backward transfer.
+However, WMAR has a slight advantage over WMAR (FIFO only), where almost no backward transfer is observed; see Table 2.
+Further, WMAR shows more consistent performance across tasks with fewer fluctuations in performance during training, which may be a desirable property if there is a need to take a model snapshot and use the agent at arbitrary points during training.
 Forward Transfer
-Our results show no evidence of inhibition to the strength of the agent when it is required to learn multiple tasks with shared structure.
-Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer, see Table 1.
-We find significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance.
+There is no evidence of inhibition to the strength of the agent when learning multiple tasks.
+Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer; see Table 2.
+We found significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance.
 Importantly, the ordering of the tasks is such that levels in later tasks are either subsets of, or different to levels in earlier tasks.
-This is so that the agent is unable to completely generalise on earlier tasks just by learning later tasks.
+This is so that the agent cannot completely generalise on earlier tasks just by learning later tasks.
 Crucially, our results show that strong forward transfer also holds for tasks that are not subsets of earlier tasks.
-Figure 3: Continual learning performance of WMAR with (top) or without (bottom) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained.
+Figure 4: Continual learning performance on tasks with shared structure.
+Bold line segments denote the periods in which certain tasks are trained.
 Scores are normalised by Equation 1.
@@
 $0.32\pm 0.06$
-4.2 Tasks without Shared Structure
-We use a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and follow Machado et al.
-(2018) in using sticky actions in Atari environments.
-We run the CL experiment on an arbitrary ordering of the environments as testing each permutation would be prohibitively expensive.
-We show that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL.
-Moreover, lacking a task oracle, we find predetermined reward scales (Table S.LABEL:S-tab:atari_rew_scales) for each individual environment necessary to enable continual learning.
-Baseline Performance
-WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Section S.LABEL:S-sec:baseline for parameters and results).
-Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
-Continual Learning
-The normalised performance is plotted in Figure 4 (tabular results in Section LABEL:S-sec:cl).
-It is clear that DreamerV3 is not capable of CL; it only performs well when it is actively training on a given task.
-Forgetting
-WMAR shows significantly improved stability in learning at all points after the task is learned.
-With only the FIFO buffer, and in all tasks except the last, the agent loses virtually all performance before the following task has finished training.
-This is in direct contrast to WMAR with the addition of the long-term global distribution matching buffer where instead, the agent is able to consistently maintain much of its performance on tasks learned earlier on (Table 2).
-This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model.
-As training progresses, samples from the first task have a decreasing probability of being drawn.
-This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn.
-Despite this, our results show that WMAR is able to hold much of its performance on the earliest of tasks (Figure 4).
-Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
-Figure 4: Continual learning performance of WMAR with (top) or without (middle) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained.
-DreamerV3 benchmark results are also given (bottom).
+5 Discussion
+WMAR has the ability to perform CL with reduced memory requirements compared to DreamerV3.
+On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits dramatically reduced forgetting.
+However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure) remains a challenge.
+Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learnt less effectively if they are learnt later.
+Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer.
+This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers.
+Our results reiterate findings of world model approaches as strong baselines for continual RL problems, which we attribute to their natural ability to support off-policy experience replay.
+Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output.
+Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
+Generalisation of Tasks
+While we observed positive forward and backward transfer on tasks with shared structure, it was not the case for tasks without.
+We posit this is due to reliance on the convolutional feature extractor, which could benefit from commonality in the visual domain only.
+Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks.
+For example, changes such as inverting the colours and permuting the controls is likely to induce negative forward transfer.
+Reward Scaling without Task Oracle
+We found that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales.
+While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we found that this property did not hold when reward scales differed significantly (e.g.
+by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt.
+We believe that without a task oracle, this limitation affects all continual RL.
+It is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes.
+Experiments with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
+Memory Capacity
+Despite the benefits and improved memory capacity of WMAR, a key limitation of any buffer-based method is finite capacity.
+In our case, as more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting.
+Indeed, we observe some variance in the amount of forgetting to which each task is susceptible.
+Results from the experiments on tasks without shared structure (Figure 3) suggest that the latter tasks which may have been learnt poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
+6 Conclusion and Future Work
+We extended a well-known model-based ‘world model’ architecture, DreamerV3, with an improved memory buffer, and applied it to the problem of continual RL in two settings: a continuous stream of tasks with and without shared structure i.e.
+commonalities between the tasks that could be leveraged by the agents.
+Performance was evaluated for forward and backward transfer, in addition to the common practice of measuring only forgetting.
+We found that model-based agents are capable of continual learning on tasks with shared structure, with some minor benefits to the enhanced buffer in WMAR.
+And in tasks without shared structure, the improved buffer in WMAR conferred substantial improvements.
+In both cases, WMAR has substantially reduced memory requirements.
+The results suggest that model-based RL using a world model with a memory-efficient replay buffer can be an effective and practical approach to continual RL, justifying future work.
+We expect future developments to the concepts shown here, or for existing techniques like behaviour cloning in CLEAR, to be combined with our method for further performance gains.
+Such an adaptation could counter shifts in the latent distribution as the world model trains and where the actor is frozen.
+Further, we posit the possibility of tuning the hyperparameters and configuration of WMAR to increase sample efficiency and model capacity for further improvements.
+We also propose a difference in learning tasks with and without shared structure, where later tasks may be dissimilar in input, control, and reward scale, and hope for future research to explore the effect of such difference on CL within supervised and non-supervised tasks.
+Appendix A Background and Related Work
+A.1 Reinforcement Learning
+We explore the application of CL in finite, discrete time partially observable RL environments.
+Each environment may be characterised as a Partially Observable Markov Decision Process Kaelbling et al.
+(1998) (POMDP), in which the agent does not have access to the state $s\in S$, and is a generalisation of the fully observable case Puterman (1990).
+A POMDP is a tuple $M=(\mathcal{S},\mathcal{A},p,r,\Omega,\mathcal{O},\gamma)$.
+$\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions.
+Where, at time step $t$, the current state and action $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$ and the next state $s_{t+1}$ are modelled by the stochastic state transition $s_{t+1}\sim p(s_{t},a_{t})$.
+The reward is given by $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R},r(s_{t},a_{t},s_{t+1})$.
+$\Omega$ is the set of observations, where the observation $\omega_{t}\in\Omega$ is modelled by $\omega_{t}\sim\mathcal{O}(s_{t})$.
+The discount factor for computing returns is $\gamma\in(0,1)$.
+Simultaneously, actions are sampled from a stochastic policy $a_{t}\sim\pi(\omega_{t})$.
+The objective is to find a $\pi$ which maximises the expected discounted returns from any initial state $\mathbb{E}_{\pi}[R_{0}\mid s_{0}]$, where under a finite horizon $T$, $R_{t}=\sum_{i=t}^{T}\gamma^{i-t}r(s_{i},a_{i},s_{i+1})$.
+Generally, the policy $\pi_{\theta}$ is a neural network parameterised by $\theta$.
+The policy may be learnt through a model-free approach, which means that no part of the learning system exists to simulate the environment.
+RL algorithms may also be classified as on- or off-policy methods.
+On-policy methods require that new samples from the environment be generated with the latest policy $\pi_{\theta}$ for each update.
+Off-policy methods must tackle the challenge of the discrepancy between the policy’s current performance and its past performance when the sample was generated Espeholt et al.
+(2018).
+However, the ability to update the current policy using previously collected samples allows for better sample efficiency.
+An alternative is model-based approaches with learned world models.
+Here, the RL algorithm learns a simulator to model the state transitions $p$ and rewards $r$ using information from observations from past rollouts.
+Then, an RL algorithm may act in the simulator to learn the policy.
+Further, samples collected from any stage of learning may be stored in a replay buffer and used to train the simulator, making this method off-policy.
+This method involving the proxy of learning a simulator has also been shown to enable superior sample efficiency.
+A.2 Continual Learning
+Although we exclusively explore the application and development of CL techniques in the RL setting, the problem of CL has also long existed in the supervised learning domain, attracting a wide body of research (e.g.
+Kirkpatrick et al.
+(2017); Lopez-Paz & Ranzato (2017)).
+Neural networks have long been known to exhibit the catastrophic forgetting phenomenon.
+Indeed, the problem was first recognised within the context of supervised learning with neural networks McCloskey & Cohen (1989); Chen & Liu (2018), where it was discovered that training on new tasks or categories will likely cause a neural network to forget the learnings from previous tasks.
+In the common supervised learning setting, a dataset may be continually shuffled and replayed, eliminating non-stationarity in the data distribution, ensuring balanced performance across all samples French (1999).
+This approach has been shown to be an adequate solution to catastrophic forgetting.
+However, it may not be possible to sample a stationary data distribution in applications where large amounts of data are streamed, such as in RL where tasks are introduced sequentially and not reliably repeated.
+A.3 Continual Reinforcement Learning
+In time, the challenges of CL were recognised and subsequently applied to RL. Here, the challenge of CL can be seen as a special case of a supervised learning task in off-policy RL.
+Previous research has attempted to apply alternate training schemes supported by replay.
+However, these methods focus mainly on model-free RL algorithms Isele & Cosgun (2018); Rolnick et al.
+(2019).
+Generally, strategies to mitigate catastrophic forgetting can be grouped by rehearsal, regularisation, and parameter isolation methods.
+All of these methods work to reduce plasticity and forgetting and aim to stabilise learning; believing that it is stability, not plasticity, that is the key limiting factor.
+Rehearsal methods most commonly work by using replay of past data, whereas regularisation and parameter isolation methods work through constraining select parameters or the output of the neural network itself.
+Synaptic consolidation approaches include regularisation and parameter isolation methods and aim to preserve the network parameters important to previously learned tasks.
+These approaches include Elastic Weight Consolidation (EWC) Kirkpatrick et al.
+(2017) and Progress & Compress (P&C) Schwarz et al.
+(2018), which attempt to consolidate past experience at the level of individual neurons through constraining the movement of select weights.
+It is not obvious how experience replay may be used in synergy with those methods within an RL algorithm.
+Other approaches focus on replaying experiences or generating typical experiences Shin et al.
+(2017), to improve agent training and avoid catastrophic forgetting.
+These methods have achieved strong results.
+The state-of-the-art CL algorithm in model-free RL is the CLEAR method Rolnick et al.
+(2019), which uses replay buffers augmented with V-trace importance sampling for off-policy learning and behaviour cloning to improve learning stability.
+Inspired in part by evidence from neuroscience, alternate augmentations to the replay buffer, such as selective experience replay, favouring surprise or reward, have also been proposed.
+Empirical data Isele & Cosgun (2018) from experiments examining the efficacy of various replay buffer augmentations, determined that matching the global training distribution to be an effective method of mitigating catastrophic forgetting without the requirement to store all experiences.
+World models Ha & Schmidhuber (2018) describe model-based RL approaches that are able to solve a variety of complex environments with superior sample efficiency Hafner et al.
+(2023; 2022); Schrittwieser et al.
+(2020).
+Research Kessler et al.
+(2022) related to the applicability of world models to continual learning has shown evidence of their potential in this domain.
+The work has also shown that the DreamerV2 Hafner et al.
+(2022) algorithm with a standard FIFO replay buffer that persists across tasks was able to curb catastrophic forgetting.
+Such an approach may be thought of as a memory approach.
+However, typical approaches to training world models and continual learning with world models demand large replay buffers, holding on the order of millions of samples, most notably in the form of images with a significant memory footprint.
+This poses a computational constraint and challenge to scalability.
+Rather, we are able to show that, through augmentation of the replay buffer, a significantly smaller, specifically crafted buffer can also curb catastrophic forgetting.
+Appendix B World Models
+B.1 Training Algorithm
+The overall training algorithm is shown in Algorithm 1.
+Algorithm 1 World Models with Augmented Replay
+Hyperparameters: World model training iterations $K$.
+Input: World model $M$, augmented replay buffer $\mathcal{D}$, sequence of tasks $\boldsymbol{\mathcal{T}}_{1:T}=(\tau_{1},\tau_{2},\dots,\tau_{T})$.
+for $\tau=\tau_{1},\tau_{2},\dots,\tau_{T}$ do
+     for $i=1,2,\dots,K$ do
+         Train world model $M$ on $\mathcal{D}$.
+         Train actor $\pi$ using $M$.
+         Use $\pi$ in $\tau$ and append episodes to $\mathcal{D}$.
+     end for
+end for
+B.2 Network Architecture
+The Actor Critic architecture is shown in Figure 5.
+Figure 5: Actor critic definition.
+We adhered to most of the parameters and architectural choices of DreamerV3.
+We found that the listed differences had little impact on agent performance.
+Changes were primarily made to benefit wall time, as running continual learning experiments is computationally expensive.
+See Table 3.
+Table 3: Summary of differences in hyperparameters.
+Name DreamerV3 WMAR WMAR (FIFO only)
+Replay capacity (FIFO) 1M 0.26M 0.52M
+Replay capacity (long-term) 0 0.26M 0
+Batch size 16 16 16
+Batch length 64 32 32
+Activation LayerNorm+SiLU LayerNorm+SiLU LayerNorm+SiLU
+Activation (GRU) LayerNorm+SiLU Tanh Tanh
+CNN Encoder and Decoder
+Following DreamerV3, the design of the convolutional feature extractor takes as input a $64\times 64$ RGB image as a resized environment frame.
+The encoder convolutional neural network (CNN) LeCun et al.
+(1989) consists of stride 2 convolutions of doubling depth with the same padding until the image is at a resolution of $4\times 4$, where it is flattened.
+We elect to use the “small” configuration of the hyperparameters controlling network architecture from DreamerV3 to appropriately manage experiment wall time.
+Hence, 4 convolutional layers are used with depths of 32, 64, 128, and 256 respectively.
+As with DreamerV3 we use channel-wise layer normalisation Ba et al.
+(2016) and SiLU Hendrycks & Gimpel (2023) activation for the CNN.
+The CNN decoder performs a linear transformation of the model state to a $4\times 4\times 256=4096$ vector before reshaping to a $4\times 4$ image and inverting the encoder architecture to reconstruct the original environment frame.
+MLP
+All multi-layer perceptrons (MLP) within the RSSM, actor, and critic implement MLPs with 2 layers and 512 hidden units in accordance with the “small” configuration of DreamerV3.
+B.3 Replay Buffer
+Algorithm 2 Sampling from combined buffers $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$
+Combined (augmented) buffer $\mathcal{D}\dot{=}\{\mathcal{D}_{1},\mathcal{D}_{2}\}$.
+Uniformly sample $i\in\{1,2\}$.
+return Sampled minibatch from $\mathcal{D}_{i}$.
+B.4 Task Agnostic Exploration
+Previous research uses the Plan2Explore algorithm Sekar et al.
+(2020) to facilitate exploration within different tasks due to a lack of a task oracle.
+Plan2Explore is a natural extension to the Dreamer algorithm, working by training an ensemble of predictors which each attempt to model the following state’s convolutional feature extractor output.
+The variance of the predictions is then used as a measure of the world model’s uncertainty, thus suggesting actions that lead to states that are good for exploration.
+Appendix C Experiments
+C.1 Procgen Environments
+We used the CoinRun game in Procgen. The CoinRun environments are defined in Table 4.
+Table 4: Procgen CoinRun environment names and configuration.
+Name No backgrounds Restricted themes Monochrome assets
+CoinRun - - -
+CoinRun+NB X - -
+CoinRun+NB+RT X X -
+CoinRun+NB+RT+MA X X X
+C.2 Experimental Parameters and Execution Time
+Our method makes use of a number of environment samples significantly larger than the size of the replay buffer.
+Constraints on computational resources necessitate approximate conversions of GPU time.
+We follow Hafner et al.
+(2023) by estimating that the A40 is twice as fast as the V100.
+We also estimate that the A40 is marginally faster than the L4.
+The environment steps and samples, and the compute time in approximate GPU days are multiplied by the number of tasks in the CL suite.
+Appendix D Results
+D.1 Validation of Implementation
+To further validate that our implementation is a faithful representation of DreamerV3, we compared WMAR (FIFO only) to an open source implementation by the author https://github.com/danijar/dreamerv3.
+We tested it on the tasks without shared structure.
+See Figure 6 and compare to Figure 3.
+Figure 6: Continual learning performance of DreamerV3, with bold line segments denoting the periods in which certain tasks are being trained.
 Scores are normalised by Equation 1.
 The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
-Table 2: Average forgetting and forward transfer on tasks without shared structure.
-Model Avg. forgetting Avg. fwd. transfer
-(Lower is better)
-(Higher is better)
-WMAR
-$\mathbf{0.10}\pm 0.04$
-$-0.35\pm 0.10$
-WMAR (FIFO)
-$0.68\pm 0.10$
-$\mathbf{-0.11}\pm 0.14$
-Forward Transfer and Plasticity
-We find significant differences in the properties of continual learning on tasks with and without shared structure which are reflected in the results.
-Contrary to findings in tasks with shared structure, the continual learning setting offers no benefit to forward transfer.
-Benefits to continual learning in greatly reduced forgetting come at the cost of also reduced forward transfer — effectively trading some plasticity for a significant amount of stability, as the agent is unable to learn later tasks as effectively.
-We observe a decreasing trend in the agent’s normalized score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’ (Table S.LABEL:S-tab:raw_res_atari), which is already well below 1.0.
-This evidences that it performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the agent.
-This limitation thus results in plasticity being gradually lost as training progresses.
-Moreover, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learned a previous task as compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in CL with world models.
-5 Discussion
-WMAR has the ability to perform CL with reduced memory requirements.
-On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits reduced forgetting.
-However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure), remains a challenge.
-Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learned less effectively if they are learned later.
-Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer.
-This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers.
-Further, being able to learn tasks quickly benefits scalability of the algorithm.
-Our results reiterate findings of world model approaches as strong baselines for CL problems, which we attribute to their natural ability to support off-policy experience replay.
-Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output.
-Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
-5.1 Generalisation of Tasks
-While we find tasks with shared structure to enable positive forward and backward transfer, we note that this is not the case in tasks without shared structure.
-We posit this is due to the reliance on the convolutional feature extractor within the WMAR architecture, enabling it to only take advantage of commonality in the visual domain.
-Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks.
-For example, a change to an environment involving inverting the colours and permuting the controls is likely to induce negative forward transfer.
-A key distinction between CL with world models and model-free methods is the requirement to reconstruct the high-dimensional observations in the form of images.
-We posit this induces significant strain on the capacity of the model, possibly in the learning of the latent space, or in the model parameters for encoding and reconstructing the observation, especially if it varies greatly between tasks.
-5.2 Reward Scaling without Task Oracle
-We also find that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales.
-While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we find this property to not hold when reward scales differ significantly (e.g.
-by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt.
-Barring the possibility of access to a task oracle, we believe that this limitation affects all continual learning applications.
-This limitation is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes.
-Experiments involving approaches with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
-5.3 Limitations of Reservoir Sampling
-The WMAR approach brings substantial improvement to memory requirements and CL characteristics, all while significantly reducing catastrophic forgetting.
-However, the scalability of this approach to a larger number of tasks remains unproven.
-A key limitation inherited from the use of reservoir sampling is the finite capacity of the long-term global distribution matching buffer.
-As more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting.
-Indeed, we observe some variance in the amount of forgetting to which each task is susceptible.
-Results from the experiments on tasks without shared structure (Figure 4) suggest latter tasks which may have been learned poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
-A plausible approach to improving the performance of WMAR on a larger number of tasks is to increase the size of the long-term global distribution matching buffer.
-However, this approach naturally inherits limitations from world model approaches and memory-based CL approaches, as doing so incurs significant memory requirements.
-While techniques such as image compression may alleviate this constraint, RL is frequently done at scale, where the need to send and store a significant volume of image data may pose a significant engineering challenge, reducing the applicability and accessibility of this algorithm.
-6 Conclusion and Future Work
-We develop a powerful approach for CL within the context of RL approaches using world models.
-We expect future developments to the concepts shown here, or for existing techniques like behaviour cloning in CLEAR to be combined with our method for further performance gains.
-Such an adaptation could be countering shifts in the latent distribution as the world model trains and where the actor is frozen.
-Further, we posit the possibility of tuning the hyperparameters and configuration of WMAR to increase sample efficiency and model capacity to further improve CL.
-We also propose a difference in learning tasks with and without shared structure, where later tasks may be dissimilar in input, control, and reward scale, and hope for future research to explore the effect of such difference on CL within supervised and non-supervised tasks.
+D.2 Baseline – single task runs
+The parameters used for single tasks for CoinRun (shared structure) are shown in Table 5 and for Atari (without shared structure) in Table 6.
+Table 5: CoinRun single task training parameters. GPU days are measured in A40 days.
+Env. steps Env. samples Approx. GPU days Buffer size
+PPO (CNN-LSTM) 4M 4M 0.25 On policy
+DreamerV3 1.6M 1.6M 0.25 1M
+WMAR 1.5M 1.5M 0.25 0.26M $+$ 0.26M
+WMAR (FIFO only) 1.5M 1.5M 0.25 0.52M
+Table 6: Atari single task training parameters. GPU days are measured in A40 days.
+Env. steps Env. samples Approx. GPU days Buffer size
+PPO (CNN-LSTM) 10M 2.5M 0.25 On policy
+DreamerV3 100K 0.4M 0.1M 0.25 1M
+WMAR 6M 1.5M 0.25 0.26M $+$ 0.26M
+WMAR (FIFO only) 6M 1.5M 0.25 0.52M
+The single task results for CoinRun (shared structure) are shown in Table 7 and for Atari (no shared structure) in Table 9 with the reward scales in Table 8
+Table 7: Single task Procgen CoinRun experimental results across 5 random seeds. Scores are unnormalized scores at the end of training ($p^{(\tau)}_{\tau}(N)$).
+Task Random PPO (CNN-LSTM) DreamerV3 WMAR
+CoinRun 2.7 6.0 6.5 6.6
+CoinRun+NB 2.7 6.0 9.0 7.0
+CoinRun+NB+RT 2.7 6.1 8.8 7.0
+CoinRun+NB+RT+MA 2.7 6.0 9.3 7.5
+Table 8: Atari environment reward scales for training WMAR.
+Name Reward scale
+Ms Pacman 0.05
+Boxing 1
+Crazy Climber 0.001
+Frostbite 0.2
+Table 9: Single task Atari experimental results across 5 random seeds. Scores are unnormalized scores at the end of training ($p^{(\tau)}_{\tau}(N)$).
+Task Random PPO (CNN-LSTM) DreamerV3 100K WMAR
+Ms Pacman 12 1631 1375 1737
+Boxing 0 87 76 89
+Crazy Climber 8 103939 87706 116097
+Frostbite 15 274 921 836
+D.3 Continual Learning
+On the Continual Learning tasks, the results for CoinRun are shown in Table 10 and for Atari in Table 11.
+Table 10: Normalised results for each CoinRun environment.
+All metrics are average and standard errors over 5 seeds.
+EOS denotes the End Of Segment (i.e.
+where the bold segment ends in Figure 4 in the main document), and EOT denotes the End Of Training.
+Environment WMAR WMAR (FIFO only)
+EOS EOT EOS EOT
+CoinRun $0.93\pm 0.06$ $1.01\pm 0.04$ $0.89\pm 0.05$ $0.86\pm 0.08$
+CoinRun+NB $1.06\pm 0.03$ $1.23\pm 0.02$ $0.96\pm 0.05$ $1.10\pm 0.04$
+CoinRun+NB+RT $1.16\pm 0.03$ $1.21\pm 0.03$ $1.19\pm 0.03$ $1.22\pm 0.03$
+CoinRun+NB+RT+MA $1.09\pm 0.03$ $1.09\pm 0.03$ $1.18\pm 0.02$ $1.18\pm 0.02$
+Table 11: Normalised results for each Atari environment.
+All metrics are average and standard errors over 5 seeds.
+EOS denotes the End Of Segment (i.e.
+where the bold segment ends in Figure 3 in the main document), and EOT denotes the End Of Training.
+Environment WMAR WMAR (FIFO only) DreamerV3
+EOS EOT EOS EOT EOS EOT
+Ms Pacman $1.08\pm 0.13$ $0.91\pm 0.07$ $1.06\pm 0.10$ $0.11\pm 0.06$ $0.73\pm 0.02$ $0.10\pm 0.01$
+Boxing $0.84\pm 0.04$ $0.86\pm 0.03$ $1.01\pm 0.04$ $-0.05\pm 0.02$ $0.91\pm 0.01$ $0.09\pm 0.01$
+Crazy Climber $0.45\pm 0.09$ $0.19\pm 0.06$ $0.67\pm 0.03$ $-0.05\pm 0.01$ $0.77\pm 0.02$ $0.01\pm 0.00$
+Frostbite $0.62\pm 0.32$ $0.62\pm 0.32$ $2.42\pm 0.64$ $2.42\pm 0.64$ $3.47\pm 0.52$ $3.47\pm 0.52$